Four litres of 0.4M 3-(N-morpholino) propane sulfonic acid (MOPS) buffer of pH 7.5 are required. Available in the laboratory are 10M HCl and 10M NaOH and the zwitterionic form of MOPS (M.Wt = 210.3). The pKa for the MOPS zwitter ion is 7.2. Describe, showing the relevant calculations, the procedure for preparing the buffer 


1
Expert's answer
2021-09-15T02:22:01-0400

Molar mass of 209.2633 g/mol

209.2633 × 0.4 = 83.71g

4000/83.71 = 47.78 moles

10 × 10/47.78

= 100/47.78 =2.09

7.5+2.09 = 9.09